# ABScribe: Rapid Exploration of Multiple Writing Variations in Human-AI Co-Writing Tasks using Large Language Models

### v1

@@ 5. Results
“The user interface for A [ABScribe] made comparisons, easy storage and access of those variations much easier than B [Baseline]. The fact that after you wrote a prompt, it instantly assigned a button to it that you could access later, was incredibly useful. It made it such that you could actually play with a more precise number of variations than I previously could and the fact that you could manually edit them and then again, quickly, have a way to play with the variation made it much more practical as a writing tool, there was a lot less physical effort involved to streamline that process. It was a wonderful, very smart way of dealing with the problem of clutter on the page.” – W1
The overall response to ABScribe, as exemplfied by by W1’s comment and Figures 7, 7 and 8, was positive, with a significant increase () in the summed agreement levels on the efficacy of the revision process (RQ1), and a significant reduction () in TLX rating for subjective task workload (RQ2) when compared to the Baseline interface.
To gain deeper insights into the factors behind the reduction in task workload and the increase in user perceptions, we conducted a reflexive thematic analysis on the semi-structured user interviews. User perceptions on the revision process groups findings relating to how the users perceived changes to the process and outcome of creating and managing variations as well as their interactions with the LLM. Subjective task workload groups findings relating to the ease of specific tasks during the writing and revision process as well the ease of specific interactions with the interface to accomplish those tasks.
@@ 5.1. RQ1: User Perceptions on the AI-Assisted Revision Process
F1: ABScribe lessens pressure to commit early to an initial idea and nudges users to explore a greater quantity of variations than the Baseline workflow. The non-linear approach to storing multiple variations within AB Text Components without cluttering the document, and the ability to switch between variations using the Hover Buttons and the Variation Accordion made some participants feel less pressured to commit to an initial variation before considering multiple options. For example, W3 noted: “I would probably feel more pressured to just kind of work on one sentence and come up with a couple of variations and change it immediately. I would feel like I have to commit pretty early on, rather than generating a number of variations, trying a bunch of different tracks and sort of different timelines, seeing how each of them turns out and performing a master comparison at the end. So I think it had a significant effect, or would have a significant effect on my behavior, certainly doing the same task for both conditions where I was trying to deal with a bunch of different versions and trying to change them and revise them differently. It was vastly more difficult in B [Baseline] and a nonlinear approach seems to make much more sense.”
@@
W2,3 and 12 specifically commented that the lack of interface clutter due to the way AB Components store variations and the ability to reuse prompts using the automatically generated Variation Buttons meant that they could create, revise and evaluate a larger number of variations with ABScribe. W12 noted that storing variations in a linear flow is “just very clunky”, leading to a “higher cost of keeping multiple versions”, “more scrolling in the document”, and “taking a longer time to find anything specific”. While exploring variations with Baseline, W12 noted that: “I sort of gave up about halfway because it was taking too long to vaguely remember, oh there was this version that I thought was good, but I wasn’t able to find it because the document became so long, so I just grabbed whatever to just finish.”
F2: ABScribe enhances writers’ ability to explore more granular variations in context of the surrounding passage. W1, 3, and 9 mentioned that they could work with smaller text-segments when using ABScribe more easily than Baseline. When using the chat-based LLM interface in Baseline, W3 said they had “never even considered [editing] on a sentence level because it would be so hard to go into ChatGPT and say that in the third sentence of the third paragraph…I don’t even know if it [ChatGPT] has a sense of where that is.” This sentiment was echoed by W9 who noted: ‘‘Usually, [I] would be [editing] at least a whole paragraph and see edits from that and then paste some of those edits in [to the document] but I wouldn’t put one sentence in…so there isn’t very fine-grained control [in Baseline]…whereas here, [with ABScribe], because it takes less effort, I’m like okay, I can do one sentence.”
@@
Being able to work with smaller text-segments, however, made some participants worry about the overall coherence of the passage. W9 said “I fear that if I edit small chunks, then the tone of different chunks end up becoming different. Whereas I kind of want to change the tone of the whole thing to one specific thing.” W3 shared a similar concern but noted that the ability to view and edit variations in-place, within the context of the surrounding passage reduced their concern: “ABScribe is vastly superior for any kind of fine-grained edits which become incredibly difficult to deal with [in Baseline] if you’re trying to do different edits on smaller variations of the text, unless you want to perform a single edit and immediate make that change.”
@@
F3: ABScribe nudges writers toward an imperative LLM prompt writing style in contrast to the conversational style in Baseline. W1, 3, 6 and 8 noted how the automatically generated Variation Buttons that captured prompts they previously wrote influenced their prompt writing style to be more direct. W1 mentioned that in ABScribe, “you’re conscious of the fact that you are designing a button, and that forces you think within that framework. You’re not really asking someone to do something [like in the chat-based interface]. You’re giving it instructions to make a button…which make things very simple, very easy, very quick…making it much easier to create prompts, and then use them subsequently, when they’re being instantly converted to a new use-case.” The conversational approach sometimes led users to anthropomorphize the AI, using superfluous words that do not influence the quality of the variation generated. For example, we observed that W8 would say “Can you please make this variation shorter?” when instructing the chatbot, but opt for a more direct “make it shorter” prompt for ABScribe. When asked why, they said the conversational AI assistant felt “similar to Clippy”, referencing the Office Assistant from a discontinued intelligent interface from Microsoft with an interactive animated character (Baym et al., 2019).
@@
F4: ABScribe nudges writers toward composing generalizable and atomic LLM prompts in contrast to the complex and variation-specific prompts in the conversational approach in Baseline. W3 and 12 shared that with ABScribe, they were intentionally trying to simplify their prompt design to make them generalizable and reusable across different text segments. W3 said “I felt more like I was going to create a generalized prompt, that I will probably reuse later. So it felt like it ought to be something that’s simple that could be applied to a variety of situations rather than something that’s specific to a single piece of text.” In contrast to W3, W9 found that due to the ease of prompt reuse, they were more likely to create longer prompts that would generalize to other contexts, when compared to the chat interface. “I was okay with writing longer prompts, for example, to imitate the style of a character because it took less effort, and it was fun for me to do, and I knew I could use it again in other sentences. Whereas in [Baseline], I wasn’t as excited because it would take more time to type, and I knew that if I had to reuse it, I would need to type it again.”
@@
However, because participants were actively thinking of instructions they could reuse across different text segments, they were less likely to write prompts that were tied to specific characteristics of particular text-segments, as exemplified by W3’s comment: “I categorically preferred A [ABScribe] to B [Baseline]…the only advantage to B might be that it slightly encouraged me to have more nuance in the prompts I gave to ChatGPT…it just made me realize I could do that. I know I could do that in ABScribe too, but the button interface kind of guided me to more naturally consider simplistic prompts rather than prompts which might be more suited to particular tasks for those in conversation with a bot.” Decomposing custom prompts that only apply to a specific text segment into more general, more atomic prompts that apply across variations encouraged participants to combine multiple prompts by clicking multiple Variation Buttons. W3 considered mimicking the functionality of longer, more complex prompts by “stacking” multiple of their general, more atomic prompts to explore variations: “I think I would also be more willing to try out a variety of permutations of different prompts, rather than trying to apply one prompt globally to an entire email, or full prompts to one or two sentences. So seeing how compound prompts might help at various points would be a lot better in A [ABScribe].”
@@ 5.2. RQ2: Subjective Task Workload
Participants pointed to four aspects of the revision workload in ABScribe that led to the significant reduction in subjective task workload: reduction in clutter, ease of variation management, access to surrounding text context, and reduced context-switching during LLM use.
@@
F5: ABScribe reduces task workload by reducing document clutter. W4, 5, 7, 10, 11, commented that the general lack of clutter in the ABScribe interface when dealing with multiple versions affected various aspects of their writing and revision process. Specifically, the way ABScribe stores variations in-place with the ABToolbar as opposed to in-sequence in the doucument reduced clutter. W11 summarizes this in the following quote “I think definitely the biggest difference would have been the fact that your your document is not as cluttered. Actually, it doesn’t get cluttered basically because you can just switch between versions and the text is on in the same location. So that’s already a huge boost in terms of not having a mess on my hand. And you notice at the very start I was already organized to think that way.” When comparing ABScribe to the linear storage of variations in Baseline, W5 and 8 noted that the linear interface inevitably led to cluttered and messy documents. W5 pointed out the difficulties of managing variations in the linear Baseline approach: “So the linear approach was more difficult, because there was just walls of text, like they began piling up very quickly. And I tried to segment them, right, I believe, if I numbered them, it would have been better. But at the same time, it doesn’t get rid of the root problem, where more and more text is, is being added to the whole draft.”
@@
Although all participants agreed that the ABScribe interface produced a less cluttered, in contrast to what we expected, two of them perceived clutter as being not necessarily bad. W2 and W6, who were both very comfortable with the worklow afforded by Baseline as it was similar to what they were used to doing, noted that clutter didn’t matter as much to them during the revision process. W2 mentioned that Baseline “definitely” felt more familiar, and that “…you probably think cluttering is one of the one of the important factors to consider when people are writing, I actually don’t think that’s the case. That’s why I don’t care about whether it looks [messy during revision].” W6, who described themselves as a “a messy editor”, a “hoarder” of various made copies of older text, felt “really conflicted” about the reduction in clutter because they liked being able to “mishmash multiple versions together” in a messy document. In describing the workflow in Baseline, they said it was “more familiar to me than like doing it the way that you would in ABScribe. Even though in an abstract way, the non-linear approach makes a lot of sense…I feel like it just feels like that’s how the design should be…but I feel like [I’m] a messy editor. And so it’s, it’s almost easier for me to edit, in Baseline.”
@@
F6: ABScribe reduces task workload by enhancing variation management. Two major recurring activities in the exploration of multiple variations were variation storage, or tracking variation history, as W7 referred to it, and the comparison of multiple variations. W3, 6, 7, and 11 noted that the non-linear storage of the ABScribe interface necessitated less overhead to manage and revise multiple variations. W3 commented that by having the ABScribe interface manage storing variations for them, they were more able to focus on the writing task: “You’re fully focused on the writing changes you’re trying to make, as opposed to managing the state and managing your document and managing like, that kind of stuff. So that was the biggest difference for me. So I really liked that feature. And that made a huge difference in general to the task. But because I’m less focused on management of things, or management of my thoughts a little bit, it’s a lot easier just using that, like the versioning system.”
@@
While W3 touched on the ease of variation storage, W11 discussed how the Hover Buttons enhanced the ease with which they compared variations: “the feature made it easier to to do the comparisons, because then you can click the version that you’re - you’re comparing with, and then hover and look at the text [for the other variations]. Whereas, with the Baseline, you have to both keep track of where the version you’re comparing with is and also simultaneously figure out which version you’re comparing to…so that is trickier than dealing with the new approach.”
Some users noted that there were some instances where comparing larger text-segments was easier to do in-sequence, and suggested that the Variation Accordion interface could also serve the purpose of in-sequence comparison. W3 notes this: “Yeah, I mean, one nice thing about Baseline is that I do get to see all of the versions together. So they’re all listed for me…but it’s much harder to have a number of sentences, which you’re generating different versions for because even once you start to hit two different sentences, [and] we’re trying to generate different variations, the document becomes very cluttered, and becomes difficult to manage. And you forget what the context is for each of those different variations. So for - for context, and for clutter, I think A is vastly superior.”
F7: ABScribe reduces task workload by showing variations in context of the surrounding text during manual editing and LLM use. The need for considering context during the revision process came up during several interviews. W3-8, and 10 commented that thanks to the in-place comparison of variations, they were able to see what a variation looks like in a paragraphs, as exemplified by W10’s comment: “I like the nonlinear version, because when you hover on the different buttons, you can directly see the impact of different variations within the paragraph or within the context. So in that way, you know, whether the text fit into the original document or not. Whereas in Baseline, if you put all the variations linearly in the document, at some point, you just start to lose a sense of what’s the context of this of this sentence, what am I writing there? Also, the chatbot in Baseline, I’d say it’s pretty much [the] same as ChatGPT. So if I want something, I need to scroll back to try to look for it. So that’s pretty much similar to the current AI writing system.”
@@
W7 echoes this preference for the non-linear ABScribe interface: “When you’re writing a paragraph, you’re not looking at a sentence in isolation. So if you’re changing a particular sentence, you want to see how it looks in comparison to the rest of your text. And so to have the nonlinear version allows you to kind of do that more seamlessly than with a linear version, where you’d have to reorganize a lot more in order to have that effect.” Whereas, W10 mirrored this, and noted that manually organizing variations while simultaneously figuring out context was challenging: “I need to manually think of a way to organize all the variations so that I understand what they mean. Or like, what, how they’re connected to the original text. That cost a lot of time. And it’s like, very high physical demand.”
@@
F8: ABScribe’s variation storage and in-place LLM revision reduces subjective task workload . Interactions that brings the user outside of the primary text editing interface and break their flow of writing or revising were perceived to be effortful and time-consuming. W7 points out how the baseline interface leads to these sorts of interactions: “So if I were to go into version history, and I wanted to go back to a very particular change in one particular paragraph, but I made that like 50 changes ago, I would either have to revert back to something where all of the document would have been unchanged, or I’d have to do like a very inconvenient and kind of cumbersome process of like copying that particular change from that particular version history into my current doc and then proceeding, which is, like tedious” W4, 5, 7, and 9 commented that they all had various ways of managing different variations of text when using the baseline or outside of the study that required them to leave the primary text editing interface to perform comparison of variations. This was either to copy paste different versions from separate documents into the primary text editing interface for comparison in sequence, or simply to compare variations side by side in separate documents. W5, 7, and 9 all found the ABScribe interface less cumbersome, especially when performing edits on several smaller text segments, due to the lack of context-switching. W9 explains this here: “Whereas here [ABScribe], I think because it takes less effort, like okay, I can do one sentence. I also want to do another one. So I’ll do that. I don’t need to copy the whole paragraph in [the chat-interface of Baseline] and try to get an answer from that. I can just do it to those sentences.”
@@
F9: ABScribe reduces task workload by making prompts more reusable than Baseline. Almost all users (W2, 3, 6-10, 12) noted that prompt reuse was much easier in the ABScribe interface. W3 commented on how the baseline interface imposed a “memory load or cognitive load issue to remember what prompts you have.” W2 notes how the AI buttons of the ABScribe interface eliminate the need to rewrite prompts stating “I don’t need to rewrite prompts every time. It was really very quick and efficient. The usability of this one in terms of buttons, the reusing the prompts is very good.” While W8 called ABScribe’s AI buttons “much more streamlined.” W3 sums up the interaction concisely in the following quote: “A [ABScribe] is vastly superior for reusability, there is no question. B [Baseline], you basically have to work from memory, which can also be fine. But with version A, you click it, you don’t have the same memory load or cognitive load issue to remember what prompts you have before, cannot be more better facilitated.”

### v3

@@ 5. Results
“The user interface for A [ABScribe] made comparisons, easy storage and access of those variations much easier than B [Baseline]. The fact that after you wrote a prompt, it instantly assigned a button to it that you could access later, was incredibly useful. It made it such that you could actually play with a more precise number of variations than I previously could and the fact that you could manually edit them and then again, quickly, have a way to play with the variation made it much more practical as a writing tool, there was a lot less physical effort involved to streamline that process. It was a wonderful, very smart way of dealing with the problem of clutter on the page. Even absent the AI generation part of it’s an excellent way of storing variations.” – W1
The overall response to ABScribe, as exemplfied by by W1’s comment and Figures 7, 7 and 8, was positive, with a significant increase () in the summed agreement levels on the efficacy of the revision process (RQ1), and a significant reduction () in TLX rating for subjective task workload (RQ2) when compared to the Baseline interface. To gain deeper insights into the factors behind the reduction in task workload and the increase in user perceptions, we conducted a reflexive thematic analysis on the semi-structured user interviews. User perceptions on the revision process (section 5.1) groups findings relating to how the users perceived changes to the process and outcome of creating and managing variations as well as their interactions with the LLM. Subjective task workload (section 5.2) groups findings relating to the ease of specific tasks during the writing and revision process as well the ease of specific interactions with the interface to accomplish those tasks.
@@ 5.1. RQ1: User Perceptions on the AI-Assisted Revision Process
ABScribe provides features targeted at improving Variation Management and thoughtful use of AI in text revision. To disentangle results of the two as well as identify when findings are a result of both those features, findings in the following section are tagged. We use Variation Management, AI Integration, or Mixed to indicate whether a finding relates purely to the management of text variations, or how participants interact with the AI, or a mixture of both features. When a finding is Mixed, comments are provided to exemplify what users attributed to Variation Management elements, and what was due to AI Integration elements.
@@
F1: Variation Management–ABScribe lessens pressure to commit early to an initial idea and nudges users to explore a greater quantity of variations than the Baseline workflow. The non-linear approach to variation storage within a Variation Field and the ability to interact with them via the Popup Toolbar and Variation Sidebar without cluttering the document made some participants feel less pressured to commit to an initial variation before considering multiple options. In contrast, participants indicated that in traditional text editing settings, they maintained and iterated on a single variation. During the study, we noticed baseline interactions favoured expansion of variations for a single passage, revision, then collapse to a single variation to avoid clutter, whereas with ABScribe, multiple variations were maintained for considerations. For example, W3 notes how the difficulty of variation storage in Baseline pushed them to maintain just one variation: “I would probably feel more pressured to just kind of work on one sentence and come up with a couple of variations and change it immediately. I would feel like I have to commit pretty early on, rather than generating a number of variations, trying a bunch of different tracks and sort of different timelines, seeing how each of them turns out and performing a master comparison at the end. So I think it had a significant effect, or would have a significant effect on my behavior, certainly doing the same task for both conditions where I was trying to deal with a bunch of different versions and trying to change them and revise them differently. It was vastly more difficult in B and a nonlinear approach seems to make much more sense.”
@@
W2, 3, and 12 specifically commented that the lack of interface clutter due to the way the Variation Fields, Popup Toolbar, and Variation Sidebar handle variation storage meant that they could create, revise and evaluate a larger number of variations with ABScribe. W12 noted that variation storage in a linear flow is “just very clunky”, leading to a “higher cost of keeping multiple versions”, “more scrolling in the document”, and “taking a longer time to find anything specific”. While exploring variations with Baseline, W12 noted that: “I sort of gave up about halfway because it was taking too long to vaguely remember, oh there was this version that I thought was good, but I wasn’t able to find it because the document became so long, so I just grabbed whatever to just finish.”
F2: Mixed–ABScribe enhances writers’ ability to explore more granular variations in context of the surrounding passage. W1, 3, and 9 mentioned that they could work with smaller text-segments more easily when using ABScribe than Baseline. Participants indicated that this was due to both the Popup Toolbar enabling in-context comparison of variations (meaning they could switch between variations in-place observing how they would fit in to the passage) and the AI Modifiers allowing for less effortful use of AI on finer-grained text segments.
@@
With regards to variation management, a concern participants presented when editing fine-grained text segments within traditional text editing interfaces is that the overall coherence of the passage would be more difficult to maintain while performing variation comparison. W9 said “I fear that if I edit small chunks, then the tone of different chunks end up becoming different. Whereas I kind of want to change the tone of the whole thing to one specific thing.” W3 shared a similar concern but noted that the ability to view and edit variations within the context of the surrounding passage reduced their concern: “ABScribe is vastly superior for any kind of fine-grained edits which become incredibly difficult to deal with [in Baseline] if you’re trying to do different edits on smaller variations of the text, unless you want to perform a single edit and immediate make that change.”
With regards to AI Integration, When using the chat-based LLM interface in Baseline, W3 said they had “never even considered [editing] on a sentence level because it would be so hard to go into ChatGPT and say that in the third sentence of the third paragraph…I don’t even know if it [ChatGPT] has a sense of where that is.” This sentiment was echoed by W9 who noted: ‘‘Usually, [I] would be [editing] at least a whole paragraph and see edits from that and then paste some of those edits in [to the document] but I wouldn’t put one sentence in…so there isn’t very fine-grained control [in Baseline]…whereas here, [with ABScribe], because it takes less effort, I’m like okay, I can do one sentence.”
The combination of AI Modifiers and Popup Toolbar enhances writers’ ability to explore more granular variations by enabling the writer to remain within-context as they edit.
@@
F3: AI Integration–ABScribe nudges writers toward an imperative LLM prompt writing style in contrast to the conversational style in Baseline. W1, 3, 6 and 8 noted how the automatically generated AI Modifers that captured prompts they previously wrote influenced their prompt writing style to be more direct. W1 mentioned that in ABScribe, “you’re conscious of the fact that you are designing a button, and that forces you think within that framework. You’re not really asking someone to do something [like in the chat-based interface]. You’re giving it instructions to make a button…which make things very simple, very easy, very quick…making it much easier to create prompts, and then use them subsequently, when they’re being instantly converted to a new use-case.” Some examples of the short imperative style from the participants include “Make more professional, Make more casual, write this so a fifth grader could understand it” Note the imperative style and general nature of the prompts.
In contrast, the conversational approach sometimes led users to anthropomorphize the AI, using superfluous words that do not influence the quality of the variation generated. For example, we observed that W8 would say “Can you please make this variation shorter?” or “Could you make a version with more varied vocabulary”” when instructing the chatbot, but opt for a more direct “make it shorter” prompt for ABScribe. When asked why, they said the conversational AI assistant felt “similar to Clippy”, referencing the Office Assistant from a discontinued intelligent interface from Microsoft with an interactive animated character (Baym et al., 2019).
@@
F4: AI Integration–ABScribe nudges writers toward composing generalizable and atomic LLM prompts in contrast to the complex and variation-specific prompts in the conversational approach in Baseline. W3 and 12 shared that with ABScribe, they were intentionally trying to simplify their prompt design to make them generalizable and reusable across different text segments. W3 said “I felt more like I was going to create a generalized prompt, that I will probably reuse later. So it felt like it ought to be something that’s simple that could be applied to a variety of situations rather than something that’s specific to a single piece of text.” In contrast to W3, W9 found that due to the ease of prompt reuse, they were more likely to create longer prompts that would generalize to other contexts, when compared to the chat interface. “I was okay with writing longer prompts, for example, to imitate the style of a character because it took less effort, and it was fun for me to do, and I knew I could use it again in other sentences. Whereas in [Baseline], I wasn’t as excited because it would take more time to type, and I knew that if I had to reuse it, I would need to type it again.”
@@
However, because participants were actively thinking of instructions they could reuse across different text segments, they were less likely to write prompts that were tied to specific characteristics of particular text segments, as exemplified by W3’s comment: “I categorically preferred A [ABScribe] to B [Baseline]…the only advantage to B might be that it slightly encouraged me to have more nuance in the prompts I gave to ChatGPT…it just made me realize I could do that. I know I could do that in ABScribe too, but the button interface kind of guided me to more naturally consider simplistic prompts rather than prompts which might be more suited to particular tasks for those in conversation with a bot.” A prompt pattern in the conversational approach was to copy paste a section of text then ask for a modification such as the following by W3 ‘‘”Comprehending the essence of crisp, contemporary material …” Can you write this in a way that maximized the number of linkedin likes?”
With ABScribe, participants tended to decompose specific prompts that only apply to a given text segment into more general, atomic prompts that apply across variations. To achieve similar specificity, participants combined multiple prompts by clicking selecting multiple AI Modifiers. W3 considered mimicking the functionality of longer, more complex prompts by “stacking” multiple of their general, more atomic prompts to explore variations: “I think I would also be more willing to try out a variety of permutations of different prompts, rather than trying to apply one prompt globally to an entire email, or full prompts to one or two sentences. So seeing how compound prompts might help at various points would be a lot better in A [ABScribe].” W8 demonstrates this stacking of atomic prompts in the following use of the AI Drafter “@ai tell me a joke, make it more creative, personalize!, add a cough or a speech utterance between every few words” Note the use of several generic and atomic prompts ”make it more creative” followed by ”personalize!”
@@ 5.2. RQ2: Subjective Task Workload
Participants pointed to four aspects of the revision workload in ABScribe that led to the significant reduction in subjective task workload: reduction in clutter, ease of variation management which referenced both variation storage and variation comparison, access to surrounding text context, and the reduction of context switching during LLM use.
@@
F5: Variation Management–ABScribe reduces task workload by reducing document clutter. W4, 5, 7, 10, 11, commented that the general lack of clutter in the ABScribe interface when dealing with multiple versions affected various aspects of their writing and revision process. Specifically, the way ABScribe handles variation storage in-place (i.e., all variations of a segment of text are stored and can be accessed at the Variation Field using the Popup Toolbar) as opposed to in-sequence, by storing both versions in the main text editing interface, reduced clutter. W11 summarizes this in the following quote “I think definitely the biggest difference would have been the fact that your your document is not as cluttered. Actually, it doesn’t get cluttered basically because you can just switch between versions and the text is on in the same location. So that’s already a huge boost in terms of not having a mess on my hand. And you notice at the very start I was already organized to think that way.” When comparing ABScribe to the linear storage of variations (one after the other within the main document) in Baseline, W5 and 8 noted that the linear interface inevitably led to cluttered and messy documents. W5 pointed out the difficulties of managing variations in the linear Baseline approach: “So the linear approach was more difficult, because there was just walls of text, like they began piling up very quickly. And I tried to segment them, right, I believe, if I numbered them, it would have been better. But at the same time, it doesn’t get rid of the root problem, where more and more text is, is being added to the whole draft.”
A noteworthy and surprising trade-off is that although all participants agreed that the ABScribe interface produced less clutter, two of them perceived clutter as unproblematic or even beneficial. W2 and W6, were both comfortable with the workflow afforded by Baseline as it was similar to what they were used to doing in their naturalistic writing tasks - paper, article, and book writing. They noted that clutter didn’t matter as much to them during the revision process. W2 mentioned that Baseline “definitely” felt more familiar, and that “…you probably think cluttering is one of the one of the important factors to consider when people are writing, I actually don’t think that’s the case. That’s why I don’t care about whether it looks [messy during revision].” W2 also mentioned when doing variation comparison: “In terms of comparison, I think this is just, I don’t know, it just, it looks convenient. But actually, when you’re actually writing things that matter you don’t want to just use this to compare the sentences. … For example, I guess most people use ChatGPT to write emails, that’s the most popular application, when they write emails to clients and professors, people—they don’t want to offend them. But sometimes you want to pay attention to your wording and stuff, in that case, I don’t think anyone wants to just compare like this. You want to be careful.” and concluded with “I would not compare them on this page [via the Popup Toolbar and hover feature], I would compare them in the Idea Bucket [Variation Sidebar].”
@@
W6, who described themselves as a “a messy editor”, a “hoarder” of various made copies of older text, felt “really conflicted” about the reduction in clutter. They commented: ‘‘I feel like it would be a little bit easier to do that [variation comparison], because the text is already there [in the main document editor] … you can’t see two versions side by side unless you were to scroll in the box [Variation Sidebar], right. And so I feel like version B would be just like, it’s easier there to do it. [variation comparison]” and that they liked being able to “mishmash multiple versions together” in a messy document. In describing the workflow in Baseline, they said it was “more familiar to me than like doing it the way that you would in ABScribe. Even though in an abstract way, the non-linear approach makes a lot of sense…I feel like it just feels like that’s how the design should be…but I feel like [I’m] a messy editor. And so it’s, it’s almost easier for me to edit, in Baseline.”
@@
F6: Variation Management–ABScribe reduces task workload by enhancing variation management. Two major activities in the exploration of multiple variations were variation storage, or tracking variation history, as W7 referred to it, and variation comparison. W3, 6, 7, and 11 noted that ABScribe’s non-linear storage of variations necessitated less overhead to track variation history and do variation comparisons . W3 commented that by having the ABScribe interface manage variation storage for them as opposed to manually placing text variations in the document, they were more able to focus on the writing task: “You’re fully focused on the writing changes you’re trying to make, as opposed to managing the state and managing your document and managing like, that kind of stuff. So that was the biggest difference for me. So I really liked that feature. And that made a huge difference in general to the task. But because I’m less focused on management of things, or management of my thoughts a little bit, it’s a lot easier just using that, like the versioning system.”
@@
While W3 touched on the ease of variation storage, W11 discussed how the Popup Toolbar enhanced the ease with which they did variation comparison: “the feature made it easier to to do the comparisons, because then you can click the version that you’re - you’re comparing with, and then hover and look at the text [for the other variations]. Whereas, with the Baseline, you have to both keep track of where the version you’re comparing with is and also simultaneously figure out which version you’re comparing to…so that is trickier than dealing with the new approach.”
A trade-off some users commented on was that it was easier to compare large segments of text in-sequence within the main document as is done in the Baseline. W3 and W7 both noted the benefit for comparing variations in-sequence within the document. W3 said “Yeah, I mean, one nice thing about Baseline is that I do get to see all of the versions together [viewed at the same time in the main text editing space]. So they’re all listed for me” and W7 stated “I think in that particular scenario[taking a small part of variation 1 and merging it with most of variation 2 in order to make my best variation], having a linear approach [like Baseline] may be more helpful for me to like, see all the different variants and then compare?” However, they both followed their statements by noting the Variation Sidebar could serve a similar purpose so the trade-off was perceived as advantageous for ABScribe. W3 commented: “[with Baseline] it’s much harder to have a number of sentences, which you’re generating different versions for because even once you start to hit two different sentences, [and] we’re trying to generate different variations, the document becomes very cluttered, and becomes difficult to manage. And you forget what the context is for each of those different variations. So for context, and for clutter, I think A is vastly superior.” and W7 said “your idea bucket [Variation Sidebar] does afford us that feature as well. Because I found myself doing that a lot. I was checking when, when we were first testing out the tool and giving me comps, I was checking to see if it was different enough. And we weren’t I wasn’t constantly like toggling in the document text, Right? So I was just like scanning the idea buckets [Variation Sidebar], I don’t think that you’re losing out on that feature, necessarily, just because I’m seeing that the linear approach is more helpful”
@@
F7: Variation Management–ABScribe reduces task workload by showing variations in context of the surrounding text during manual editing and when editing with AI. The need for considering context during the revision process came up during several interviews. The benefits of variation management on editing variations with context are exemplified by W3-8, and 10’s comments. They note that thanks to the in-placevariation comparison via the hover interaction with the Popup Toolbar, they were able to see what a variation looks like within a paragraph. To this effect, W10’s states: “I like the nonlinear version, because when you hover on the different buttons, you can directly see the impact of different variations within the paragraph or within the context. So in that way, you know, whether the text fit into the original document or not. Whereas in Baseline, if you put all the variations linearly in the document, at some point, you just start to lose a sense of what’s the context of this of this sentence, what am I writing there?” W7 echoes the value of viewing variations with surrounding context: “When you’re writing a paragraph, you’re not looking at a sentence in isolation. So if you’re changing a particular sentence, you want to see how it looks in comparison to the rest of your text. And so to have the nonlinear version allows you to kind of do that more seamlessly than with a linear version, where you’d have to reorganize a lot more in order to have that effect.” Whereas, W10 mirrored this, and noted that manually organizing variations while simultaneously figuring out context was challenging: “I need to manually think of a way to organize all the variations so that I understand what they mean. Or like, what, how they’re connected to the original text. That cost a lot of time. And it’s like, very high physical demand.”
@@
F8: Mixed–ABScribe’s variation storage and in-place LLM revision reduces subjective task workload. Interactions that bring the user outside of the primary text editing interface and break their flow of writing or revising were perceived to be effortful and time-consuming. W7 points out how variation management–particularly variation storage - with the Baseline interface necessitates context switching: “So if I were to go into version history, and I wanted to go back to a very particular change in one particular paragraph, but I made that like 50 changes ago, I would either have to revert back to something where all of the document would have been unchanged, or I’d have to do like a very inconvenient and kind of cumbersome process of like copying that particular change from that particular version history into my current doc and then proceeding, which is, like tedious”
W4, 5, 7, and 9 commented that they all had various ways of handling variation management in their naturalistic writing tasks such as producing articles, research papers, and books, which were similar to how they handled variations with the Baseline interface. These methods, such as copy-pasting different versions from separate documents into the primary text editing interface for comparison in-sequence, or doing variation comparison side by side in separate windows required them to leave the primary text editing interface to perform variation comparison. W5, 7, and 9 all found variation comparison less cumbersome with ABScribe, especially when performing edits on several smaller text segments, due to the lack of context switching. W9 explains this here: “Whereas here [ABScribe], I think because it takes less effort, like okay, I can do one sentence. I also want to do another one. So I’ll do that. I don’t need to copy the whole paragraph in [the chat-interface of Baseline] and try to get an answer from that. I can just do it to those sentences.”
Use of AI within the Baseline was also indicated as a source of context switching. W10 notes how the chat-like interface in Baseline pulls them out of context: ‘‘Also, the chatbot in Baseline, I’d say it’s pretty much [the] same as ChatGPT. So if I want something, I need to scroll back to try to look for it. So that’s pretty much similar to the current AI writing system.” ABScribe’s AI modifiers and AI Drafter enable the use of AI without requiring leaving the main text editing interface.
@@
F9: AI Integration– ABScribe reduces task workload by making prompts more reusable than Baseline. Almost all users (W2, 3, 6-10, 12) noted that prompt reuse was much easier in the ABScribe interface. W3 commented on how the Baseline interface imposed a “memory load or cognitive load issue to remember what prompts you have.” W2 notes how the AI Modifiers of the ABScribe interface eliminate the need to rewrite prompts stating “I don’t need to rewrite prompts every time. It was really very quick and efficient. The usability of this one in terms of buttons, the reusing the prompts is very good.” While W8 called ABScribe’s AI Modifiers “much more streamlined.” W3 sums up the interaction concisely in the following quote: “A is vastly superior for reusability, there is no question. B, you basically have to work from memory, which can also be fine. But with version A, you click it, you don’t have the same memory load or cognitive load issue to remember what prompts you have before, cannot be more better facilitated.”